OpenStudy (anonymous):

Algebra 2: Which equation has the set of real numbers as its solution? A) 2(-2r+4)= 4(-2r+6) B) 3/5 (fraction) (z+4)= 4/5 (Fraction) z C) 2x+5(x-3) -6=-8+7x-13 D) 0.7-(q+3) = 3.3-q SHOW WORK TO ANSWER

OpenStudy (turingtest):

the answer is all reals if the equation can be reduced to 0=0 somehow

OpenStudy (turingtest):

not sure what you mean by B) 3/5 (fraction) (z+4)= 4/5 (Fraction) z what does (fraction) mean exactly ?

OpenStudy (anonymous):

\[\frac{ 3 }{ 5}(z+4)=\frac{ 5 }{ 5 }z\]

OpenStudy (turingtest):

ok, so for example lets do this one multiple both sides by 5, what do you have?

OpenStudy (anonymous):

25? Im bad at math sooo sorry if i make u mad lol

OpenStudy (turingtest):

\[\cancel5\cdot\frac3{\cancel5}=3\]

OpenStudy (anonymous):

ahhh

OpenStudy (turingtest):

so the left becomes\[3(x+4)\]what about the right side?

OpenStudy (anonymous):

5z or 1z

OpenStudy (turingtest):

\[\cancel5\cdot\frac5{\cancel5}=5\]

OpenStudy (turingtest):

so \[5z\]

OpenStudy (anonymous):

yay

OpenStudy (turingtest):

now we need to keep going\[3(z+4)=5z\]distribute to simplify the left...

OpenStudy (anonymous):

3z+12=5z

OpenStudy (turingtest):

but wait, your problem says it should be\[\frac35(z+4)=\frac45z\]is there a typo?

OpenStudy (anonymous):

nop

OpenStudy (turingtest):

well in your post you wrote B) 3/5 (fraction) (z+4)= 4/5 (Fraction) z and in the math feature you wrote\[\frac35(z+4)=\frac55z\]which is it?

OpenStudy (turingtest):

4/5 or 5/5 ?

OpenStudy (anonymous):

4/5 lol srry

OpenStudy (turingtest):

okay, so what should the right be after multiplying by 5 then?

OpenStudy (turingtest):

(everything else is still correct)

OpenStudy (anonymous):

5z+20= 5z

OpenStudy (turingtest):

no, you changed the left nothing should change on the left; you did that correctly before\ the only change is on the right...

OpenStudy (turingtest):

\[5\cdot\frac35(z+4)=5\cdot\frac45z\]simplifies to what?

OpenStudy (anonymous):

5z+20=4z idk

OpenStudy (turingtest):

you made a mistake on the left you had it correctly before\[\cancel5\cdot\frac3{\cancel5}(z+4)=\cancel5\cdot\frac4{\cancel5}z\]

OpenStudy (anonymous):

3z+12z=4z

OpenStudy (turingtest):

yes can you solve it for z?

OpenStudy (turingtest):

oh wait, you should get\[3z+12=4z\](you have an extra z in there somehow)

OpenStudy (anonymous):

ah and yea 3z+12=4z -3z -3z 12=1z z=12

OpenStudy (turingtest):

correct so there is only one answer: z=12 that is not all real numbers, just one

OpenStudy (anonymous):

so i got the answer?!?!?1 or do i do the others?

OpenStudy (turingtest):

you now know that the answer to B) is *not* all real numbers since you cannot get it to be 0=0 through algebra try A) and see what you get for that...

OpenStudy (anonymous):

on a i can use any number then if it is a real numb it is the answer

OpenStudy (turingtest):

just solve it and see what you get

OpenStudy (turingtest):

2(-2r+4)= 4(-2r+6)

OpenStudy (turingtest):

distribute the 2's on each side and you get...?

OpenStudy (turingtest):

sorry I meant distribute the 2 on the left and the 4 on the right

OpenStudy (anonymous):

-4r+8= -8r+24

OpenStudy (turingtest):

right try to solve it for r

OpenStudy (anonymous):

4r=16

OpenStudy (turingtest):

right, so r=?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

4

OpenStudy (turingtest):

right again, that is just a single answer, so the answer to A) is also not ALL REALS because again we did not get 0=0

OpenStudy (turingtest):

let's try C)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

7x^2+21-6= -8+7x-13

OpenStudy (turingtest):

not quite 2x+5(x-3) -6=-8+7x-13 distribute the 5 first

OpenStudy (anonymous):

2x+5x-15-6= -8+7x-13

OpenStudy (turingtest):

right :) now simplify

OpenStudy (anonymous):

7x-15-6= -8+7x-13

OpenStudy (turingtest):

yep now combine the constants and try to solve for x

OpenStudy (anonymous):

soo to combine 7x+7x?

OpenStudy (turingtest):

no -15-6=? -8-13=?

OpenStudy (anonymous):

7x-9=-21+7x

OpenStudy (turingtest):

-15-6 is not -9 try the left side again

OpenStudy (anonymous):

-21

OpenStudy (anonymous):

7x- -21=-21+7x +

OpenStudy (turingtest):

no you already have the - sign included... 7x+(-15-6)=( -8-13)+7x 7x+(-21)=-21+7x 7x-21=7x-21

OpenStudy (anonymous):

its the same so were done?

OpenStudy (turingtest):

right or you can notice that we can subtract 7x-21 from both sides to get 0=0 which is always true, so the answer is all reals you should still check D) to make sure that it also is not all reals, there could be more than one answer for all we know (at least from the information given it is not clear that there is only one answer)

OpenStudy (anonymous):

yay tnx <3

OpenStudy (turingtest):

welcome!